3. Find the continuous-time convolution of r(t) = 10e (i 2)u(t - 2) with h(t) = 5e ( 4)u(t – 4)
Answer #1

61W Que act = (be )(- nit= selt-4) 4(6-4) X(as) = 10 (W) e zjw HC)= (t e ti Y(W)= X() H(a) 5150 letraje cy( pestial fraction

Kindly upvote if it's helpful for you.